# Formalising the intentional stance: attributing goals and beliefs to stochastic processes

### v1

@@ Abstract
We are concerned with the behaviour of stochastic systems with inputs and outputs, and how this might relate to the pursuit of a goal. We model this using what we term transducers, which are a mathematical object that captures only the external behaviour of such a system and not its internal state. We present a framework for reasoning about the optimality of such a process, when it is coupled to a ‘teleo-environment’ consisting of another transducer that also embodies a success criterion. We find that (globally) optimal transducers have a property closely related to Bellman’s theorem: a transducer that is optimal in one time step will again be optimal in the next time step, but with respect to a different environment (obtained from the original one by a modified version of Bayesian filtering). We also consider bounded rationality and its relationship to constrained optimality, which in our framework means optimal within some subset of all transducers. We describe a condition that is sufficient for such a subset to have this Bellman property. Additionally, we show that a policy is deterministic if and only if there exists a teleo-envionment for which it is uniquely optimal among the set of all transducers; this is at least conceptually related to classical representation theorems from decision theory. This need not hold for constrained subsets; we give an example of this related to the so-called absent-minded driver problem. All of the formalism is defined using coinduction, following the style proposed by Czajka [9].
@@ 1 Introduction
In order to do this, we propose a simple mathematical framework in which systems can be described in both mechanistic and normative-epistemic terms, and the relationships between the two can be understood. We don’t aim for the most general framework possible but for the simplest one in which our main points can be expressed. We express this framework in the mathematical language of coinduction, and indeed, a secondary purpose of this paper is to demonstrate the application of coinduction to stochastic systems.
In the remainder of this section we give an overview of our framework and a description of our results, as well as a brief explanation of our coinductive approach. In Section 2 we describe relations to previous approaches, including Dennett’s work, representation theorems for expected utility theory, the free energy principle, computational mechanics and inverse reinforcement learning, as well as our own work on the ‘as-if’ approach to cognition. The mathematical content of the paper begins in Section 3.
1.1 Outline of our approach and results
In this section we give a somewhat detailed summary of the paper’s results, using a minimum of technical language.
@@ 1.1 Outline of our approach and results
In order to reason about physical systems in such a situation, we make the following assumptions: the system has a set of possible outputs or ‘actions’ $A$ and a set of possible inputs or ‘sensor values’ $S$ (both finite). Time proceeds in discrete steps and the system’s behaviour may be stochastic. We can model its externally observable behaviour in terms of a controlled stochastic process, that is, a conditional probability distribution over infinite sequences of outputs or actions $a_{1},a_{2},\dots$ given infinite sequences of inputs or sensor values $s_{1},s_{2},\dots$. This conditional distribution must have the property that the output at time $t$ is independent of all inputs received after time $t-1$. (See [20] for a similar approach.) We refer to this as the causality condition. We will see later on that controlled stochastic processes can be defined in a different way using the mathematical framework of coinduction, and we will use coinductive reasoning throughout the paper. In the context of our coinductive definition we refer to controlled stochastic processes as transducers.
@@
Figure 1: Two ways in which a theorist might describe the behaviour of a physical system. $(a)$ in a mechanistic description the theorist posits an internal state to the system such that its dynamics can explain the observable behaviour; in the case of a robot this could include hypothesised workings of internal control systems in addition to the literal mechanical gears depicted here. We formalise mechanistic descriptions in terms of stochastic Moore machines in Section 3.3.4. $(b)$ in a normative-epistemic description the theorist posits an environment that the system believes itself to be in (according to the theorist’s interpretation) and a goal that it seeks to achieve — in this case the apple — such that the system’s behaviour can be explained by acting optimally in order to reach the goal. Normative-epistemic descriptions are the focus of our paper and are formalised in terms of teleo-envionments in Section 4. We emphasise that these are two different kinds of explanation for the same behaviour of the same system, and not two different types of system or two competing hypotheses.
We assume that the theorist has access to a controlled stochastic process representing their system’s observable behaviour. This means that the theorist is not limited to observing a single sample of the system’s behaviour given a single sequence of inputs; instead they know111or at least, have a model of how the system would behave given any number of different input sequences, over any number of trials.
Given this knowledge, one question one could ask is, “given the known externally observable behaviour, what kinds of physical mechanism might exist inside the system that would give rise to its behaviour?” Although we will briefly discuss this question in Section 3.3.4 it will not be our main focus.
@@
Within this simplified model of ideal rationality, we can ask: as a system ‘evolves’ its policy through interaction with its environment, how do the permissible mental state attributions change? By ‘evolving’ we mean the following: a transducer, i.e. a controlled stochastic process, can be seen as a conditional distribution of output sequences given input sequences, subject to a causality condition. However, once the transducer has emitted an output and received an input we can obtain a new distribution over output sequences given input sequences, by conditioning on the input-output pair and shifting both sequences by one time step, as described in sections 3.1 and 3.2.
@@
We show in section 5 that, if $\pi$ is optimal for $\varepsilon$, then after ‘evolution’ with a sensorimotor pair $(s,a)$ the evolved policy $\pi^{\prime}$ is optimal for a similar evolution $\varepsilon^{\prime}$ of $\varepsilon$. We argue in that section that this can be seen as a form of Bayesian filtering. In other words, among the many beliefs that are permissibly attributable to the system after it receives new information, one is guaranteed to look like an update of the previous attributed belief according to Bayes’ rule. (There is some subtlety about exactly what data is conditioned on, however, which will be discussed shortly.)
@@
We propose a general intuition along the following lines. Consider any notion of optimality for a control policy $x$ in an environment $y$, and some notion of updating over time for the policy and environment (we will be deliberately vague about what this ‘updating’ means in general). Suppose something like Bellman’s principle holds, namely that if a policy $x$ is optimal for $y$, then the evolution $x^{\prime}$ by a single time step will be optimal for the evolution $y^{\prime}$ of $y$. Again, $y$ will function as a permissible ‘mental state’ attribution for $x$ (in the sense that no inconsistency arises from supposing that $x$ is an ideal agent trying to solve the problem represented by $y$) and similarly for $y^{\prime}$ and $x^{\prime}$. Consequently, the ‘update’ operator that we apply to $y$ to obtain $y^{\prime}$ can be seen as updating beliefs attributed to $x$, to obtain beliefs that can be attributed to $x^{\prime}$, given any information recieved by $x$ during the update. In this sense the ‘update’ rule for $y$ can be seen as a model of idealised rationality, which generalises Bayes’ rule in some sense.
@@
When we work through the details for our teleo-environment model, we obtain an interesting insight. One might reasonably guess that a Bayesian model of rational mental state update might emerge from our model. It does, but with an important nuance: so to speak, the system is legitimated to update its ‘belief’ in a Bayesian manner, conditioning on the sensorimotor signals observed in the previous time step, provided that it additionally conditions on not having achieved its goal in the previous time step. That is, it conditions on its reieved sensor value $s\in S$, but it also conditions on $\bot$, the lack-of-success signal, even though it doesn’t actually have access to the success signal, and indeed the true value of the success signal might have been $\top$. We call this ‘value-laden filtering’, and it arises from the relevant Bellman-like principle (in Theorem 5.0.2).
@@
This is a little surprising, since an external theorist, observing the sensorimotor interactions between the system and its environment (without any additional information about whether a goal had been achieved), and applying Bayesian reasoning themselves to predict future interactions, would condition only on the sensorimotor signals. In this way, the properties of the optimality criterion in our model mean that we can derive a permissible rule for pragmatically optimal ‘mental state’ change (namely, value-laden filtering) which differs subtly from the standard Bayesian picture, whilst still being recognisable as Bayesian filtering. In future, it may be of interest to explore further how different pragmatic notions of optimality can induce different permissible update rules for mental state, and how these rules relate to ‘purely normative’ models of ideal mental state update.
@@
Another question one might want to ask is whether a given behaviour can be specified by a given teleo-environment. That is, for a system with given behaviour, does there exist a teleo-environment for which it is uniquely optimal? We explore this question in Section 6. Our results here are as one might expect from a framework based on utility-maximisation. Firstly, a behaviour can only be uniquely optimal for a given teleo-environment if it is deterministic, since otherwise it would be a mixture of two or more deterministic behaviours, each of which would also be optimal. Secondly, the specifiable behaviours are exactly the deterministic ones: if the behaviour is deterministic then there exists a teleo-environment for which it is unqiuely optimal. (There may be many such teleo-environments, but in particular there is always one in which the system must perform exactly the specified behaviour, otherwise it immediately permanently loses all hope of achieving success.) This is shown in Theorem 6.3.2. Thus, at least when it comes to deterministic behaviour, there is no difference in the set of behaviours that can be specified in a normative-epistemic way and the set of behaviours that can be specified by a framework based on internal state.
@@
This is in a sense a ‘no-go theorem’. One might initially have the intuition that systems that are ‘agents’ (in the sense of being optimal for some teleo-environment) might have special features to their behaviour, which would then allow one to test the hypothesis that some system is an agent by observing its behaviour. (See [31], which takes a Bayesian approach along these lines.) Although Theorem 6.3.2 is not a mathematically surprising result it does suggest a challenge that such an approach would need to overcome.

### v2

# Formalising the intentional stance 1: attributing goals and beliefs to stochastic processes

@@ Abstract
This article presents a formalism inspired by Dennett’s notion of the intentional stance. Whereas Dennett’s treatment of these concepts is informal, we aim to provide a more formal analogue. We introduce a framework based on stochastic processes with inputs and outputs, in which we can talk precisely about interpreting systems as having normative-epistemic states, which combine belief-like and desire-like features. Our framework is based on optimality but nevertheless allows us to model some forms of bounded cognition.
One might expect that the systems that can be described in normative-epistemic terms would be some special subset of all systems, but we show that this is not the case: every system admits a (possibly trivial) normative-epistemic interpretation, and those that can be uniquely specified by a normative-epistemic description are exactly the deterministic ones. Finally, we show that there is a suitable notion of Bayesian updating for normative-epistemic states, which we call value-laden filtering, since it involves both normative and epistemic elements. For unbounded cognition it is always permissible to attribute beliefs that update in this way. This is not always the case for bounded cognition, but we give a sufficient condition under which it is.
This paper gives an overview of our framework aimed at cognitive scientists, with a formal mathematical treatment given in a companion paper.
@@ 1 Introduction
In order to do this, we propose a simple mathematical framework in which systems can be described in both mechanistic and normative-epistemic terms, and the relationships between the two can be understood. We don’t aim for the most general framework possible but for the simplest one in which our main points can be expressed. We describe the framework intuitively in this paper, with the technical details present in the companion work.
The current work follows a particular conceptual methodology, closely related to Dennett’s intentional stance; we describe this methodology, and its relations to Dennett’s work in Section 2. Section 3 explains the mathematical framework we are proposing, and summarises the mathematical results described in the companion paper. In Section 4 we describe relations to previous formalisms: namely, representation theorems for expected utility theory, the free energy principle, computational mechanics and inverse reinforcement learning. Finally, we discuss limitations of our framework and possible future work in Section 5.
2 The ‘As-If’ Approach And Dennett’s Intentional Stance
This idea can be seen as an attempt to formalise Dennett’s intentional stance [7, 9, 10]; in doing so, we are also able to formally model relationships between (analogues of) Dennett’s intentional stance and his physical stance. The physical stance relates to technical concepts of stochastic Moore machines and unifilar machines, described in the companion paper, while the intentional stance corresponds to our exploration of the relationship between transducers and teleo-environment, in Section 3.2 and throughout both papers.
In the past there have been significant disagreement about what sorts of system should be regarded as ‘agents’. Some authors (e.g. [27, p.34] or [32]) categorise even simple feedback systems (such as thermostats or guided torpedoes) as agents. Others (e.g. [11] or [17]) explicitly reject such a characterisation, claiming that thermostats may behave (in some ways) as if they were agents, but are not truly agents. A discussion of this general debate can be found in [19]. Currently, echoes of this disagreement can be seen in disputes about the extent to which AI systems such as large language models (LLMs) should be understood as agents. The idea of ‘as-if’ agency is that we should not ignore simple systems such as thermostats when studying agency, even if the skeptics are correct and thermostats lack some important property. It is still valuable to have a theory that explains in what sense even a thermostat behaves as though it were an agent.
We will not define what a ‘formal theory of rationality’ is, but our notion of teleo-environment (Section 3.2) will serve as an example.
In the current paper and its companion, we are also concerned with consistency. We show that we obtain a form of consistency (Section 3.3) if certain assumptions are met. The form of consistency is not what we might naively expect, since the update rule is given by ‘value-laden filtering’ (Theorem Theorem), rather than vanilla Bayesian filtering. We also observe that this ‘value-laden filtering’ consistency does not hold in general in bounded-rational contexts. In particular, in Section 3.5, we describe a case in which it doesn’t hold, due to constraints on memory, and mention another consequence of optimality under memory constraints, namely that the uniquely optimal policy may be non-deterministic, as previously observed in the absent-minded driver problem ([25]) and discussed more generally in [16].
3 The Teleo-Environment Framework
In this section we describe our mathematical framework intuitively, and provide a detailed summary of our conceptually relevant results, using a minimum of technical language. We encourage the reader interested in technical details to read the companion paper, which provides a rigorous treatment.
@@ 3 The Teleo-Environment Framework
3.1 Transducers
In order to reason about physical systems in such a situation, we make the following assumptions: the system has a set of possible outputs or ‘actions’ $A$ and a set of possible inputs or ‘sensor values’ $S$ (both finite). Time proceeds in discrete steps and the system’s behaviour may be stochastic. Classically, we would model its externally observable behaviour in terms of a controlled stochastic process, that is, a conditional probability distribution over infinite sequences of outputs or actions $a_{1},a_{2},\dots$ given infinite sequences of inputs or sensor values $s_{1},s_{2},\dots$. This conditional distribution must have the property that the output at time $t$ is independent of all inputs received after time $t-1$. (See [15] for a similar approach.) We refer to this as the causality condition.
However, we choose instead to model it using an object we call a transducer, which could be described as follows.
Definition (3.0.1 in the companion paper).
Given finite sets of inputs $I$ and outputs $O$ a transducer (from $I$ to $O$) is a mathematical object that provides a distribution over $O$ and deterministically changes into another transducer when provided with a sensory value and an action chosen from its distribution.
If we start with a transducer $\pi$, give it an input $i$ and observe that it produced the output $o$, we obtain a new transducer that we refer to as “$\pi$ evolved by $(i,o)$.” Intuitively a transducer represents a process that simultaneously receives some input data and stochastically produces an output, after which it is ready for another step, perhaps in a different state.
This definition may seem not be well founded, since we are defining transducers in terms of themselves without a base case. However, a mathematical framework known as coinduction allows such definitions to be treated rigorously. The companion paper uses coinduction to reason about transducers defined in this way, and briefly explores the relation between transducers and controlled stochastic processes.
We assume that the theorist has access to a policy representing their system’s observable behaviour. This is simply a transducer from $S$ to $A$. This means that the theorist is not limited to observing a single sample of the system’s behaviour given a single sequence of inputs; instead they know111or at least, have a model of how the system would behave given any number of different input sequences, over any number of trials.
3.2 Teleo-environments
@@
Figure 1: Two ways in which a theorist might describe the behaviour of a physical system. $(a)$ in a mechanistic description the theorist posits an internal state to the system such that its dynamics can explain the observable behaviour; in the case of a robot this could include hypothesised workings of internal control systems in addition to the literal mechanical gears depicted here. We formalise mechanistic descriptions in terms of stochastic Moore machines in the companion paper. $(b)$ in a normative-epistemic description the theorist posits an environment that the system believes itself to be in (according to the theorist’s interpretation) and a goal that it seeks to achieve — in this case the apple — such that the system’s behaviour can be explained by acting optimally in order to reach the goal. Normative-epistemic descriptions are the focus of our paper and are formalised in terms of teleo-environments. We emphasise that these are two different kinds of explanation for the same behaviour of the same system, and not two different types of system or two competing hypotheses.
Given knowledge of a system’s observable behaviour in the form of a policy, one question one could ask is, “given the known behaviour, what kinds of physical mechanism might exist inside the system that would give rise to its behaviour?” Although we briefly discuss this question in Section 3.3.4 of the companion paper, it will not be our main focus.
@@
3.3 Filtering and the Bellman property
Within this simplified model of ideal rationality, we can ask: as a system ‘evolves’ its policy through interaction with its environment, how do the permissible mental state attributions change?
@@ 3.3 Filtering and the Bellman property
We show that, if $\pi$ is optimal for $\varepsilon$, then after evolution with a sensorimotor pair $(s,a)$ the evolved policy $\pi^{\prime}$ is optimal for a similar evolution $\varepsilon^{\prime}$ of $\varepsilon$. We argue in that section that this can be seen as a form of Bayesian filtering. In other words, among the many beliefs that are permissibly attributable to the system after it receives new information, one is guaranteed to look like an update of the previous attributed belief according to Bayes’ rule. (There is some subtlety about exactly what data is conditioned on, however, which will be discussed shortly.)
@@
We propose a general intuition along the following lines. Consider any notion of optimality for a control policy $x$ in an environment $y$, and some notion of updating over time for the policy and environment (we will be deliberately vague about what this ‘updating’ means in general). Suppose something like Bellman’s principle holds, namely that if a policy $x$ is optimal for $y$, then the evolution $x^{\prime}$ by a single time step will be optimal for the evolution $y^{\prime}$ of $y$. Again, $y$ will function as a permissible ‘mental state’ attribution for $x$ (in the sense that no inconsistency arises from supposing that $x$ is an ideal agent trying to solve the problem represented by $y$) and similarly for $y^{\prime}$ and $x^{\prime}$. Consequently, the ‘update’ operator that we apply to $y$ to obtain $y^{\prime}$ can be seen as updating beliefs attributed to $x$, to obtain beliefs that can be attributed to $x^{\prime}$, given any information received by $x$ during the update. In this sense the ‘update’ rule for $y$ can be seen as a model of idealised rationality, which generalises Bayes’ rule in some sense.
@@
When we work through the details for our teleo-environment model, we obtain an interesting insight. One might reasonably guess that a Bayesian model of rational mental state update might emerge from our model. In fact it does, but with an important nuance: so to speak, the system is legitimated to update its ‘belief’ in a Bayesian manner, conditioning on the sensorimotor signals observed in the previous time step, provided that it additionally conditions on not having achieved its goal in the previous time step. That is, it conditions on its received sensor value $s\in S$, but it also conditions on $\bot$, the lack-of-success signal, even though it doesn’t actually have access to the success signal, and indeed the true value of the success signal might have been $\top$. We call this ‘value-laden filtering’, and it arises from the relevant Bellman-like principle.
Theorem (Corollary 5.0.3 in the companion paper).
If a policy $\pi$ is optimal for an environment $\varepsilon$, then given a possible sequence of actions $\mathbf{a}$ and sensor values $\mathbf{s}$ the policy $\pi^{\prime}$, is optimal for the environment $\varepsilon^{\prime}$. Here, $\pi^{\prime}$ is $\pi$ evolved by $\mathbf{a}$ and $\mathbf{s}$, and $\varepsilon^{\prime}$ is $\varepsilon$ evolved by $\mathbf{a}$ and $\mathbf{s}$ (combined with a sequence of the lack-of-success symbol $\bot$).
@@
This is a little surprising, since an external theorist, observing the sensorimotor interactions between the system and its environment (without any additional information about whether a goal had been achieved), and applying Bayesian reasoning themselves to predict future interactions, would condition only on the sensorimotor signals. In this way, the properties of the optimality criterion in our model mean that we can derive a permissible rule for pragmatically optimal ‘mental state’ change (namely, value-laden filtering) which differs subtly from the standard Bayesian picture, whilst still being recognisable as Bayesian filtering.
Value-laden filtering conditions on the lack-of-success signal in addition to the received signal. Corollary 5.0.3 says that if a normative-epistemic state is a permissible attribution on one time step, then an updated version of the same normative-epistemic state will be a permissible attribution on the next time step, where the updating is given by value-laden filtering.
We are using slightly circumspect language here because of a subtle issue. Because we use “permissible attribution” to mean any teleo-environment $\varepsilon$ for which a given $\pi$ is optimal, there are in general many permissible mental state attributions for a given policy $\pi$. We might want to talk about an agent’s ‘current beliefs’ and then say that after the agent has undergone some sensor-motor interaction with its environment (i.e. been evolved by some $(\bf a,s)$), the resulting transducer $\pi^{\prime}$ has an updated set of beliefs given by $\varepsilon^{\prime}$. However, this doesn’t quite make sense, because in general there are many possible mental state attributions for $\pi^{\prime}$, so “its beliefs” are not unique. The above theorem says that the updated beliefs $\varepsilon^{\prime}$ (obtained by value-laden filtering) are always among the permissible mental state attributions for $\pi$, but there may be (and in general are) many others.
For some permissible beliefs $\varepsilon$, it happens that the updated belief $\varepsilon^{\prime}$ is the same as the one that would be obtained not by value-laden filtering but by ordinary Bayesian filtering. That is, one can obtain $\varepsilon^{\prime}$ by evolving $\varepsilon$ by $\bf a$ and $\bf s$ only, without the additional sequence of lack-of-success symbols.
In fact, given any permissible attribution $\varepsilon$, there is a way to construct another attribution, which we write $Z(\varepsilon)$, such that (i) for every policy $\pi$, the probability of achieving success at least once is equal to that for $\varepsilon$, and (ii) $Z(\varepsilon)$ has the property that $Z(\varepsilon)$-evolved-by-$\bf s$-and-$\bf a$ is a permissible attribution to $\pi^{\prime}$, i.e. we can choose to regard $Z(\varepsilon)$ as updating by ordinary Bayesian filtering instead of by value-laden filtering. The attribution $Z(\varepsilon)$ is defined in Definition 5.2.3 in the companion paper, and the proof of this statement is in Corollary 5.2.7.
To give slightly more detail, $Z(\varepsilon)$ is obtained from $\epsilon$ by a process called “single-success truncation.” It is identical to $\varepsilon$ except success will only ever occur at most once. If success does occur then $Z(\varepsilon)$ evolves into an environment that behaves identically to $\varepsilon$, except that the success signal does not occur when it otherwise would.
Both truncating environments, as well as value-laden filtering serve essentially the same purpose in modifying the update mechanism of attributed states. A transducer only represents the current state and the future of the system, completely disregarding the past. However, if an agent already achieved success in a past step, it should no longer care whether it will achieve success again. For example, consider an environment in which, on the first time step, the agent either achieves success and enters environment $X$, or doesn’t achieve success and enters environment $Y$. Environment $X$ is such that the agent must take action $A$ in order to achieve success (for the second time), while environment $Y$ is such that the agent must take action $B$. And suppose further that the agent’s sensors do not allow it to distinguish between the two environments. Even if environment $X$ is entered much more often, say with probability 0.9, an optimal agent should still ignore it and always take action $B$. The point is that in this situation, the optimal behaviour from the second time step onwards depends on the probability that success has already occurred, which is correlated with the probability of being in environment $X$ versus environment $Y$.
From the perspective of a person trying to solve this task, the reasoning at the second time step should be something along the lines of “either I am in environment $X$, in which case I can achieve success by doing action $A$, but this doesn’t matter, since I’ve already achieved success in that case, or I’m in environment $Y$, in which case I have not achieved success yet and must take action $B$ in order to achieve it at least once. Since it’s only in the case of environment $Y$ that my actions matter at all, I can discount the possibility being in environment $X$ entirely and behave as if I am in environment $Y$.”
If we were to model the agent’s knowledge as a teleo-environment but update it using vanilla Bayesian updating, it would result in an agent that always tries to achieve success starting from its current situation, regardless of whether success has already been achieved in a previous time step. Truncation avoids this by letting the agent naturally ignore any futures where it would have achieved success again, while value-laden filtering avoids it by explicitly excluding the futures that follow a success from the model.
In future, it may be of interest to explore further how different pragmatic notions of optimality can induce different permissible update rules for mental state, and how these rules relate to ‘purely normative’ models of ideal mental state update.
3.4 Specification
@@
Another question one might want to ask is whether a given behaviour can be specified by a given teleo-environment. That is, for a system with given behaviour, does there exist a teleo-environment for which it is uniquely optimal? Our results here are as one might expect from a framework based on utility-maximisation. Firstly, a behaviour can only be uniquely optimal for a given teleo-environment if it is deterministic, since otherwise it would be a mixture of two or more deterministic behaviours, each of which would also be optimal. Secondly, the specifiable behaviours are exactly the deterministic ones: if the behaviour is deterministic then there exists a teleo-environment for which it is uniquely optimal. (There may be many such teleo-environments, but in particular there is always one in which the system must perform exactly the specified behaviour, otherwise it immediately permanently loses all hope of achieving success.)
Theorem (6.2.3 and 6.3.2 in the companion paper).
Any deterministic policy $\pi$ is specifiable, i.e. there exists a teleo-environment for which $\pi$ is uniquely optimal, and furthermore any specifiable policy is deterministic.
Thus, at least when it comes to deterministic behaviour, there is no difference in the set of behaviours that can be specified in a normative-epistemic way and the set of behaviours that can be specified by a framework based on internal state.
@@
This is in a sense a ‘no-go theorem’. One might initially have the intuition that systems that are ‘agents’ (in the sense of being optimal for some teleo-environment) might have special features to their behaviour, which would then allow one to test the hypothesis that some system is an agent by observing its behaviour. (See [24], which takes a Bayesian approach along these lines.) Although Theorem Theorem is not a mathematically surprising result it does suggest a challenge that such an approach would need to overcome.
3.5 Bounded rationality